Recording Noninterest-Bearing Note Payable Entries First Choice Company buys equipment on October 1, 2020, providing as paymecCan someone let me know if this is correct?
Answer #1

Yes, It's Correct

Intrest Expenses = 2909*(3/12) = 727 (for 3 months on Dec 31,2020)